Find the curl of the function r= xi+yj over x+y


1
Expert's answer
2022-03-04T08:31:22-0500

"\text{curl }\vec F=\frac{\partial A_x}{\partial y}-\frac{\partial A_y}{\partial x}=0-0=0."